a number divided by 3 plus 6 is the same as the number minus 2, find the number.
(i know the answer is 12 i don't know how to show the work)

Answer by scott8148(6628)

let n equal the number ... (n/3)+6=n-2 ... n/3=n-8 ... n=3n-24 ... -2n=-24 ... n=12